A 0.750 kg hammer is moving horizontally at 6.50 m/s when it strikes a nail and comes to rest after driving it 1.00 cm into a board.

(a) Calculate the duration of the impact. ________S
Answer #1

Solution)

We know,

V^2=u^2+2as

So, a=-v^2/2s

Also,

V=u+at

0=u-v^2/2s*t

So, t=2s/v=2*0.01 m/6.5 m/s =0.00307 sec

=======